OpenStudy (anonymous):

If the cost of manufacturing x items is c(x) = x^3 + 21x^2 + 110x + 20 , find the marginal cost at x = 100 and the actual cost of manufacturing the 100th item.

OpenStudy (anonymous):

@electrokid can u help

OpenStudy (anonymous):

@Mertsj can u help?

OpenStudy (mertsj):

What is marginal cost? I don't know that stuff. Is the formula you gave for marginal cost?

OpenStudy (anonymous):

no its when i take the derivative of this problem then subtitute 100 in it and solve

OpenStudy (anonymous):

do you get it know?

OpenStudy (mertsj):

Did you take the derivative?

OpenStudy (anonymous):

yes its 3x^2 + 42 + 110

OpenStudy (mertsj):

Did you plug in 100?

OpenStudy (anonymous):

no hold on

OpenStudy (anonymous):

i got 90152

OpenStudy (mertsj):

\[3(100)^2+42(100)+110=3(10000)+4200+110=30000+4200=110=34310\]

OpenStudy (anonymous):

yes its 90,152

OpenStudy (mertsj):

Which equation did you plug into? The given equation or the first derivative?

OpenStudy (anonymous):

the derivative

OpenStudy (mertsj):

Well I think you need to do it again. I did it twice and got 34,310 both times

OpenStudy (anonymous):

ohh hold on

OpenStudy (mertsj):